# Optimization-free Smooth Control Barrier Function for Polygonal Collision Avoidance

### v1

@@ I Introduction
The control barrier function-based quadratic programming (CBF-QP) control method is popular for safe robotic control [1, 2, 3, 4, 5]. The CBF-based control can provide a simple and computationally efficient way for safe control synthesis [1, 2], and it has been gradually extended to higher-order systems [3, 4, 5]. Collision avoidance, i.e., driving the robot away from the obstacle and keeping a distance, is a common goal in reactive control of multi-agent robots such as [6, 7]. And recently, CBFs have been gradually used to achieve more complex collision avoidance [8, 9, 10, 11, 12, 13]. When the shapes of robots and obstacles are complicated (instead of points or spheres), collision detection is not an obvious problem. Since polytopes can non-conservatively approximate any convex shapes, compared with collision avoidance between ellipsoids or generic convex sets [8, 9, 10], references [11, 12, 13] are particularly interested in developing CBFs for collision avoidance between polytopes/polygons, where polygon refers particularly to a polytope in planar. Recently, related works about developing CBF methods for avoiding obstacles with irregular shapes can also be found in [14].
@@
This paper goes on the topic of CBFs for real-time polygonal obstacle avoidance [11, 12, 13], and inspired by but different from the approx-SDF [17], we propose an optimization-free smooth CBF, where its non-conservatism is carefully studied and guaranteed. A brief comparison with existing methods is concluded in Table I. In detail, inspired by [17], a lower bound of SDF between polygons is proposed and expressed as a nested Boolean logic composition at first, then the smooth approximation of the lower bound is established by applying the recently proposed softening method for any nested Boolean composition in [18], where the approximate error bound can be tuned by user-defined parameters. The main contributions of this paper can be summarized as follows:
@@ II-A Control Barrier Function
Definition 1 is a tiny modified version of [19, Definition 3]. The condition $h(x)<0,~{}\forall x\in\mathcal{C}^{c}$ is added to highlight that $\mathcal{C}^{c}=\left\{x:h(x)<0\right\}\neq\emptyset$, which is essential to guarantee the stability of the set $\mathcal{C}$ [1]. To provide controllers with formal safety guarantees, the control barrier function method [1] is introduced.
@@ II-B Signed Distance Field
$\operatorname{sd}(\mathcal{P}^{i},\mathcal{P}^{j})=\operatorname{sd}({0},\mathcal{P})=\small\left\{\begin{array}[]{l}~{}~{}d({0},\partial{\mathcal{P}}),{0}\notin\mathcal{P},\\ -d({0},\partial{\mathcal{P}}),{0}\in\mathcal{P}.\end{array}\right.$ (7)
@@
$\operatorname{sd}({0},\mathcal{P})=\max_{a^{\top}a=1}\min_{\omega\in\mathcal{P}}~{}a^{\top}\omega=\max_{a^{\top}a=1}-h_{\mathcal{P}}(a),$ (8)
@@ III-A Expressions of Polygons
Polygons, i.e., polytopes in planar, have a simpler geometric structure than generic polytopes. From Euler’s formula [23, Theorem 2.3.10], polygons have as many vertices as edges. Let $r^{{\mathfrak{i}}}$ denote the number of vertices and edges of $\mathcal{P}^{\mathfrak{i}}(x^{\mathfrak{i}})$ for ${\mathfrak{i}}\in\{i,j\}$. The $\mathcal{H}$)-representation is introduced to express $\mathcal{P}^{\mathfrak{i}}(x^{\mathfrak{i}})$:
@@
$A_{\mathfrak{i}}(x)$ $:={{B}\Delta v_{\mathfrak{i}}}/{\left\|\Delta v_{\mathfrak{i}}\right\|},~{}b_{\mathfrak{i}}(x):=A_{\mathfrak{i}}(x)^{\top}{v}_{\mathfrak{i}},$ (14)
@@ IV-A Establishing a lower bound of SDF
$\operatorname{sd}({0},H_{\mathcal{P}}^{-}(a))=a^{\top}0-h_{\mathcal{P}}(a)=-h_{\mathcal{P}}(a).$ (19)
@@
$h_{s}(x)=\max_{\|a\|=1}\min_{\omega\in\mathcal{P}(x)}a^{\top}\omega=\max_{a^{\top}a=1}\operatorname{sd}({0},H_{\mathcal{P}(x)}^{-}(a)),$ (20)
@@
where the set inclusion $\subseteq$” comes from the fact [24, Eq. (3.2)]: $(A\cap B)-C\subseteq(A-C)\cap(B-C)$ holds for arbitrary subsets $A,B,C$. To avoid the redundancy of symbols, sets $\mathcal{P}^{j}-\mathcal{P}^{i}_{k^{i}}$ with index ${k^{i}}$ and $\mathcal{P}^{j}_{k^{j}}-\mathcal{P}^{i}$ with index ${k^{j}}$, are together renumbered as sets $\mathcal{Q}_{k}$ with a single index $k$:
$\mathcal{Q}_{k}(x):=\left\{\begin{array}[]{l}\mathcal{P}^{j}(x^{j})-\mathcal{P}^{i}_{k^{i}}(x^{i}),k={k^{i}},~{}~{}~{}~{}~{}~{}{k^{i}}\in[r^{i}],\\ \mathcal{P}^{j}_{k^{j}}(x^{j})-\mathcal{P}^{i}(x^{i}),k=r^{i}+{k^{j}},{k^{j}}\in[r^{j}],\end{array}\right.$ (24)
@@
Recalling the fact about SDF in [25, Theorem 2.1] that: $A\subseteq B\Rightarrow\operatorname{sd}({0},A)\geq\operatorname{sd}({0},B)$, the following inequality holds naturally.
@@ IV-B Building a boolean NBF candidate from the lower bound
1) In fact, $\geq$” in (34) has been shown in Proposition 1. Hence, one only needs to prove that “$=$” in (34) holds on the mentioned point set, as proved in the following two steps.
Step 1: Recalling (31) in Lemma 2, $\mathcal{P}$ can be written as
$\mathcal{P}=\bigcap_{k\in[r^{i}+r^{j}]}\mathcal{Q}_{k}=\bigcap_{k\in[r^{i}+r^{j}]}\left(\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k}\right),$ (36)
where $\widetilde{\mathcal{Q}}_{k}=\bigcap_{k\neq m\in[r^{i}+r^{j}]}\mathcal{Q}_{m}$. Based on (36) and Lemma 4 in Appendix -B, the following relations hold:
$\partial\mathcal{P}$ $=\bigcup_{k\in[r^{i}+r^{j}]}\left(\partial\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k}\right),$ (37a)
$d({0},\partial\mathcal{P})$ $=\min_{k\in[r^{i}+r^{j}]}d({0},\partial\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k}).$ (37b)
and for every $k\in[r^{i}+r^{j}]$, one has
$\partial\mathcal{Q}_{k}$ $=(\partial\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k})\cup(\partial\mathcal{Q}_{k}/\widetilde{\mathcal{Q}}_{k}),$ (38a)
$d({0},\partial\mathcal{Q}_{k})$ $=\min\left\{d({0},\partial\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k}),d({0},\partial\mathcal{Q}_{k}/\widetilde{\mathcal{Q}}_{k}\right\}.$ (38b)
Step 2: Recalling the expressions of closed sets $\mathcal{P}$ in (6) and $\mathcal{S}$ in (16), one can obtain:
$x$ $\in{\mathcal{S}}^{c}$ $\Leftrightarrow~{}h_{s}(x)<0\Leftrightarrow$ $~{}{0}\in\operatorname{Int}(\mathcal{P}(x)),$ (39a)
$x$ $\in\partial\mathcal{S}$ $\Leftrightarrow~{}h_{s}(x)=0\Leftrightarrow$ $~{}{0}\in\partial\mathcal{P}(x),$ (39b)
$x$ $\in\operatorname{Int}(\mathcal{S})$ $\Leftrightarrow~{}h_{s}(x)>0\Leftrightarrow$ $~{}{0}\in(\mathcal{P}(x))^{c}.$ (39c)
When $x\in\partial\mathcal{S}\cup{\mathcal{S}}^{c}$, one has ${0}\in\mathcal{P}(x)$ via (39), then ${0}\in\mathcal{Q}_{k},~{}\forall k\in[r^{i}+r^{j}]$ from (36). It is obvious from ${0}\in\mathcal{P}$ that
$d({0},\partial\mathcal{P})$ $<d({0},\mathcal{P}^{*}),$ (40)
holds for any set $\mathcal{P}^{*}$ in $\mathcal{P}^{c}$. It further yields
$d({0},\partial\mathcal{P})$ $<\min_{k\in[r^{i}+r^{j}]}d({0},\partial\mathcal{Q}_{k}/\widetilde{\mathcal{Q}}_{k}).$ (41)
Applying the formulas (38b), (37b), (41) in turn, we get
$\min_{k\in[r^{i}+r^{j}]}d({0},\partial{\mathcal{Q}}_{k})$
$=\min_{k\in[r^{i}+r^{j}]}\left\{d({0},\partial\mathcal{Q}_{k}\cap\widetilde{\mathcal{Q}}_{k}),d({0},\partial\mathcal{Q}_{k}/\widetilde{\mathcal{Q}}_{k}\right\}$ (42a)
$=\min\left\{d({0},\partial\mathcal{P}),\min_{k\in[r^{i}+r^{j}]}d({0},\partial\mathcal{Q}_{k}/\widetilde{\mathcal{Q}}_{k})\right\}$ (42b)
$=d({0},\partial\mathcal{P}).$ (42c)
Recalling the facts that whenever ${0}\in\mathcal{P}$, it holds that
$\operatorname{sd}({0},\mathcal{P})$ $=-d({0},\partial\mathcal{P}),$ (43a)
$\operatorname{sd}({0},\mathcal{Q}_{k})$ $=-d({0},\partial\mathcal{Q}_{k}),~{}\forall k\in[r^{i}+r^{j}].$ (43b)
Based on this and together with (42), it yields
${h}_{a}(x)$ $=\max_{k\in[r^{i}+r^{j}]}\left\{-d({0},\partial\mathcal{Q}_{k})\right\}=-\min_{k}d({0},\partial\mathcal{Q}_{k})$
$=-d({0},\partial\mathcal{P})=\operatorname{sd}({0},\mathcal{P})={h}_{s}(x),$ (44)
which means that the inequality “$=$” in (34) has been proved.
2) First, by the aid of (39), the proposition (35) can be equivalently expressed as
${h}_{a}(x)<0$ $\Longleftrightarrow h_{s}(x)<0,$ (45a)
${h}_{a}(x)=0$ $\Longleftrightarrow h_{s}(x)=0,$ (45b)
${h}_{a}(x)>0$ $\Longleftrightarrow h_{s}(x)>0,$ (45c)
where the first proposition naturally holds from (34), and the last two are proved as follows.
Recalling (26) in Proposition 1, the left hand in (45b) becomes
${h}_{a}(x)=0\Longleftrightarrow$ ${0}\in\bigcap_{k\in[r^{i}+r^{j}]}\mathcal{Q}_{k}(x)=\mathcal{P}(x)\text{ and }$ (46)
$\exists k\in[r^{i}+r^{j}],\text{ s.t. }{0}\in\partial{\mathcal{Q}}_{k}(x),$
where the right hand in (46) is further equivalent to
[…]
In fact, by recalling (39) again, Eq. (47) is equivalent to the right hand in (45b).
[…]
$h_{s}(x)>0\Longleftrightarrow$ $~{}{0}\in\operatorname{Int}(\mathcal{P})$ (48a)
$\Longleftrightarrow$ $~{}{0}\in\operatorname{Int}({\mathcal{Q}}_{k}),~{}\forall k\in[r^{i}+r^{j}]$ (48b)
$\Longrightarrow$ ${h}_{a}(x)>0,$ (48c)
where the fact that $\operatorname{Int}(\mathcal{P})=\bigcap_{k}\operatorname{Int}({\mathcal{Q}}_{k})$ is applied. ∎
@@
First, recall a well-known property of the Linear Program (LP). From [27, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP can be obtained at the vertexes (extreme points) of the polytope (feasible domain), one knows that LPs in (27) and (28) can be reduced to
@@
2) Since ${h}_{a}(x)$ is smooth composed and due to the existing $\min/\max$ operators, according to [28, Proposition IV.9], ${h}_{a}(x)$ is piecewise $C^{1}$-continuously differentiable, which means that it is locally Lipschitz continuous. Finally, recalling that $\mathcal{S}=\mathcal{S}_{a}$ and the relationships in (35) in Theorem 1, ${h}_{a}(x)$ is a barrier function candidate on $\mathcal{S}_{a}$ by verifying Definition 1 directly. ∎
@@ V Safety Filter design with a single CBF
$\hat{\mathcal{S}}_{a}:=\hat{\mathcal{S}}_{a}({b,\kappa})=\left\{x:\hat{h}_{a}(x;\kappa,b)\geq 0\right\},$ (56)
@@
${{h}_{a}}(x)-\frac{\ln b_{2}+b}{\kappa}\leq~{}\hat{h}_{a}(x;b,\kappa)~{}\leq{{h}_{a}}(x)+\frac{\ln b_{1}-b}{\kappa},$ (57)
@@
The proof is omitted here for brevity since it is the direct corollary of [18, Theorem 5]. ∎
@@
$\nabla\hat{h}_{a}$ $=\frac{1}{\hat{\phi}+\hat{\psi}}\left(\sum_{k^{i}\in[r^{i}]}\nabla\tilde{\phi}^{k^{i}}+\sum_{k^{j}\in[r^{j}]}\nabla\tilde{\psi}^{k^{j}}\right),$ (58)
@@
$\nabla\tilde{\phi}^{k^{i}}=\frac{\nabla\hat{\phi}^{k^{i}}}{\kappa}=(\hat{\phi}^{k^{i}})^{2}\sum_{l^{j}\in[{r^{j}}]}\frac{\nabla\phi^{k^{i}}_{l^{j}}}{\hat{\phi}^{k^{i}}_{l^{j}}},$ (59a)
$\nabla\tilde{\psi}^{k^{j}}=\frac{\nabla\hat{\psi}^{k^{j}}}{\kappa}=(\hat{\psi}^{k^{j}})^{2}\sum_{k^{i}\in[{r^{i}}]}\frac{\nabla\psi^{k^{j}}_{k^{i}}}{\hat{\psi}^{k^{j}}_{k^{i}}},$ (59b)
@@
Motivated by the CBF condition (3) and the safety filter design (4), for any given locally Lipschitz continuous nominal controller $u_{0}=[(u_{0}^{i})^{\top},(u_{0}^{j})^{\top}]^{\top}$, the following distributed222 The control law (60) is distributed instead of centralized, due to that $u^{i}$ and $u^{j}$ are designed separately. QP-based control law is proposed for the input $u$ of system (15):
@@ VI-A Nonholonomic vehicle example
$\dot{p}_{x}^{\mathfrak{i}}$ $=v^{\mathfrak{i}}\cos\theta^{\mathfrak{i}},~{}\dot{p}_{y}^{\mathfrak{i}}=v^{\mathfrak{i}}\sin\theta^{\mathfrak{i}},~{}\dot{\theta}^{\mathfrak{i}}=\omega^{\mathfrak{i}},~{}\mathfrak{i}\in\{i,j\},$ (64)
@@
Here, vectors $\hat{a}$ and $\hat{\omega}^{i},\hat{\omega}^{j}$ denote the direction and points that maximize and minimize the expression in (65) respectively. These vectors depend on $x=\operatorname{col}(x^{i},x^{j})$ with $x^{i}=\operatorname{col}(p^{i}_{x},p^{i}_{y},\theta^{i})$, $x^{j}=\operatorname{col}(p^{j}_{x},p^{j}_{y},\theta^{j})$. Referring to [12, Eq. (27)], gradients $\nabla_{x^{i}}h_{s}(x),\nabla_{x^{j}}h_{s}(x)\in\mathbb{R}^{3\times 1}$ can be approximated by the local linearization:
$\nabla_{x^{i}}h_{s}(x)$ $\approx-[\hat{a}(x)^{\top},\hat{a}(x)^{\top}R(\theta^{i}+{\pi}/2)\hat{g}^{i}(x)]^{\top},$ (66a)
$\nabla_{x^{j}}h_{s}(x)$ $\approx~{}~{}[\hat{a}(x)^{\top},\hat{a}(x)^{\top}R(\theta^{j}+{\pi}/2)\hat{g}^{j}(x)]^{\top},$ (66b)
@@
Then, the SDF-linearized CBF is also simulated in Matlab/Simulink, and the comparison result is reported as follows. On the one hand, the system response under SDF-linearized CBF methods (under different sampling densities) and the proposed one are compared in Fig. 8, where the function $h_{a}(x)$ is applied to measure whether collisions happen. From the results, when the sampling density is $15$ or $10$, the SDF-linearized CBF may lead to a collision behavior, which is illustrated by the red or green lines. When the sampling density is up to $20$, the SDF-linearized CBF, as well as the proposed method, can avoid the collision between vehicles, as seen from the black and blue lines. On the other hand, the computing time of these methods can be obtained by the aid of Simulink/Profiler directly, and they are compared in Table II. From the results, one can know that the proposed method is several times faster than the SDF-linearized CBF, no matter the sampling density of the SDF-linearized CBF is chosen as $10,15$ or $20$.
As a brief summary, both methods could avoid the collision, but the proposed method is more efficient than the SDF-linearized CBF, which is a typical example of implicit, optimization-embedded methods in Table I.

### v2

@@ I Introduction
The control barrier function-based quadratic programming (CBF-QP) control method is popular for safe robotic control [1, 2, 3, 4, 5, 6, 7]. The CBF-based control can provide a simple and computationally efficient way for safe control synthesis [1, 2, 3, 4], and it has been gradually extended to higher-order systems [5, 6, 7]. Collision avoidance, i.e., driving the robot away from the obstacle and keeping a distance, is a common goal in reactive control of multi-agent robots such as [8, 9]. And recently, CBFs have been gradually used to achieve more complex collision avoidance [10, 11, 12, 13, 14, 15]. When the shapes of robots and obstacles are complicated (instead of points or spheres), collision detection is not an obvious problem. Since polytopes can non-conservatively approximate any convex shapes, compared with collision avoidance between ellipsoids or generic convex sets [10, 11, 12], references [13, 14, 15] are particularly interested in developing CBFs for collision avoidance between polytopes/polygons, where polygon refers particularly to a polytope in planar. Recently, related works about developing CBF methods for avoiding obstacles with irregular shapes can also be found in [16].
@@
This paper goes on the topic of CBFs for real-time polygonal obstacle avoidance [13, 14, 15], and inspired by but different from the approx-SDF [19], we propose an optimization-free smooth CBF, where its non-conservatism is carefully studied and guaranteed. A brief comparison with existing methods is concluded in Tab. I. In detail, inspired by [19], a lower bound of SDF between polygons is proposed and expressed as a nested Boolean logic composition at first, then the smooth approximation of the lower bound is established by applying the recently proposed softening method for any nested Boolean composition in [20], where the approximate error bound can be tuned by user-defined parameters. The main contributions of this paper can be summarized as follows:
@@ II-A Control Barrier Function
Definition 1 is a tiny modified version of [21, Definition 3]. The condition $h(x)<0,~{}\forall x\in\mathcal{C}^{c}$ is added to highlight that $\mathcal{C}^{c}=\left\{x:h(x)<0\right\}\neq\emptyset$, which is essential to guarantee the stability of the set $\mathcal{C}$ [1]. To provide controllers with formal safety guarantees, the control barrier function method [1] is introduced.
@@ II-B Signed Distance Field
$\operatorname{sd}(\mathcal{P}^{i},\mathcal{P}^{j})=\operatorname{sd}({0},\mathcal{P})=\left\{\begin{array}[]{l}~{}~{}d({0},\partial{\mathcal{P}}),{0}\notin\mathcal{P},\\ -d({0},\partial{\mathcal{P}}),{0}\in\mathcal{P}.\end{array}\right.$ (7)
@@
$\operatorname{sd}({0},\mathcal{P})=\max_{a^{\top}a=1}\min_{\omega\in\mathcal{P}}~{}a^{\top}\omega=\max_{a^{\top}a=1}-h_{\mathcal{P}}(a),$
@@ III-A Expressions of Polygons
Polygons, i.e., polytopes in planar, have a simpler geometric structure than generic polytopes. From Euler’s formula [25, Theorem 2.3.10], polygons have as many vertices as edges. Let $r^{{\mathfrak{i}}}$ denote the number of vertices and edges of $\mathcal{P}^{\mathfrak{i}}(x^{\mathfrak{i}})$ for ${\mathfrak{i}}\in\{i,j\}$. The $\mathcal{H}$)-representation is introduced to express $\mathcal{P}^{\mathfrak{i}}(x^{\mathfrak{i}})$:
@@
$A_{\mathfrak{i}}(x)$ $:={{B}\Delta v_{\mathfrak{i}}}/{\left\|\Delta v_{\mathfrak{i}}\right\|},~{}b_{\mathfrak{i}}(x):=A_{\mathfrak{i}}(x)^{\top}{v}_{\mathfrak{i}},$
@@ IV-A Establishing a lower bound of SDF
$\operatorname{sd}({0},H_{\mathcal{P}}^{-}(a))=a^{\top}0-h_{\mathcal{P}}(a)=-h_{\mathcal{P}}(a).$
@@
$h_{s}(x)=\max_{\|a\|=1}\min_{\omega\in\mathcal{P}(x)}a^{\top}\omega=\max_{a^{\top}a=1}\operatorname{sd}({0},H_{\mathcal{P}(x)}^{-}(a)),$
@@
where the set inclusion $\subseteq$” comes from the fact [26, Eq. (3.2)]: $(A\cap B)-C\subseteq(A-C)\cap(B-C)$ holds for arbitrary subsets $A,B,C$. To avoid the redundancy of symbols, sets $\mathcal{P}^{j}-\mathcal{P}^{i}_{k^{i}}$ with index ${k^{i}}$ and $\mathcal{P}^{j}_{k^{j}}-\mathcal{P}^{i}$ with index ${k^{j}}$, are together renumbered as sets $\mathcal{Q}_{k}$ with a single index $k$:
$\mathcal{Q}_{k}(x):=\left\{\begin{array}[]{l}\mathcal{P}^{j}(x^{j})-\mathcal{P}^{i}_{k^{i}}(x^{i}),k={k^{i}},~{}~{}~{}~{}~{}~{}\forall{k^{i}}\in[r^{i}],\\ \mathcal{P}^{j}_{k^{j}}(x^{j})-\mathcal{P}^{i}(x^{i}),k=r^{i}+{k^{j}},\forall{k^{j}}\in[r^{j}],\end{array}\right.$
@@
Recalling the fact about SDF in [27, Theorem 2.1] that: $A\subseteq B\Rightarrow\operatorname{sd}({0},A)\geq\operatorname{sd}({0},B)$, the following inequality holds naturally.
@@ IV-B Building a boolean NBF candidate from the lower bound
To avoid breaking the flow of the exposition, the proof is reported in Appendix -B, where the required lemma (about distance functions) is given in Appendix -C. ∎
@@
First, recall a well-known property of the Linear Program (LP). From [29, Sec 2.5, Theorem and Corollary 2] that the minimum value of a LP can be obtained at the vertexes (extreme points) of the polytope (feasible domain), one knows that LPs in (20) and (21) can be reduced to
@@
2) Since ${h}_{a}(x)$ is smooth composed and due to the existing $\min/\max$ operators, according to [30, Proposition IV.9], ${h}_{a}(x)$ is piecewise $C^{1}$-continuously differentiable, which means that it is locally Lipschitz continuous. Finally, recalling that $\mathcal{S}=\mathcal{S}_{a}$ and the relationships in (28) in Theorem 1, ${h}_{a}(x)$ is a barrier function candidate on $\mathcal{S}_{a}$ by verifying Definition 1 directly. ∎
@@ V Safety Filter design with a single CBF
$\hat{\mathcal{S}}_{a}:=\hat{\mathcal{S}}_{a}({b,\kappa})=\left\{x:\hat{h}_{a}(x;\kappa,b)\geq 0\right\},$
@@
${{h}_{a}}(x)-\frac{\ln b_{2}+b}{\kappa}\leq~{}\hat{h}_{a}(x;b,\kappa)~{}\leq{{h}_{a}}(x)+\frac{\ln b_{1}-b}{\kappa},$
@@
The proof is omitted here for brevity since it is the direct corollary of [20, Theorem 5]. ∎
@@
$\nabla\hat{h}_{a}$ $=\frac{1}{\hat{\phi}+\hat{\psi}}\left(\sum_{k^{i}\in[r^{i}]}\nabla\tilde{\phi}^{k^{i}}+\sum_{k^{j}\in[r^{j}]}\nabla\tilde{\psi}^{k^{j}}\right),$
@@
$\nabla\tilde{\phi}^{k^{i}}=\frac{\nabla\hat{\phi}^{k^{i}}}{\kappa}=(\hat{\phi}^{k^{i}})^{2}\sum_{l^{j}\in[{r^{j}}]}\frac{\nabla\phi^{k^{i}}_{l^{j}}}{\hat{\phi}^{k^{i}}_{l^{j}}},$
$\nabla\tilde{\psi}^{k^{j}}=\frac{\nabla\hat{\psi}^{k^{j}}}{\kappa}=(\hat{\psi}^{k^{j}})^{2}\sum_{k^{i}\in[{r^{i}}]}\frac{\nabla\psi^{k^{j}}_{k^{i}}}{\hat{\psi}^{k^{j}}_{k^{i}}},$
@@
Motivated by the CBF condition (3) and the safety filter design (4), for any given locally Lipschitz continuous nominal controller $u_{0}=[(u_{0}^{i})^{\top},(u_{0}^{j})^{\top}]^{\top}$, the following distributed [^2] The control law (36) is called distributed instead of centralized, due to that $u^{i}$ and $u^{j}$ are designed separately [10]. QP-based control law is proposed for input $u$ of system (13):
@@ VI-A Nonholonomic vehicle example
$\dot{p}_{x}^{\mathfrak{i}}$ $=v^{\mathfrak{i}}\cos\theta^{\mathfrak{i}},~{}\dot{p}_{y}^{\mathfrak{i}}=v^{\mathfrak{i}}\sin\theta^{\mathfrak{i}},~{}\dot{\theta}^{\mathfrak{i}}=\omega^{\mathfrak{i}},~{}\mathfrak{i}\in\{i,j\},$
@@
Here, vectors $\hat{a}$ and $\hat{\omega}^{i},\hat{\omega}^{j}$ denote the direction and points that maximize and minimize the expression in (40) respectively. These vectors depend on $x=\operatorname{col}(x^{i},x^{j})$ with $x^{i}=\operatorname{col}(p^{i}_{x},p^{i}_{y},\theta^{i})$, $x^{j}=\operatorname{col}(p^{j}_{x},p^{j}_{y},\theta^{j})$. Referring to [14, Eq. (27)], gradients $\nabla_{x^{i}}h_{s}(x),\nabla_{x^{j}}h_{s}(x)\in\mathbb{R}^{3\times 1}$ can be approximated by the local linearization:
$\nabla_{x^{i}}h_{s}(x)$ $\approx-[\hat{a}(x)^{\top},\hat{a}(x)^{\top}R(\theta^{i}+{\pi}/2)\hat{g}^{i}(x)]^{\top},$
$\nabla_{x^{j}}h_{s}(x)$ $\approx~{}~{}[\hat{a}(x)^{\top},\hat{a}(x)^{\top}R(\theta^{j}+{\pi}/2)\hat{g}^{j}(x)]^{\top},$
@@
Then, the SDF-linearized CBF is also simulated in Matlab/Simulink, and the comparison result is reported as follows. On the one hand, the system response under SDF-linearized CBF methods (under different sampling densities) and the proposed one are compared in Fig. 8, where the function $h_{a}(x)$ is applied to measure whether collisions happen. From the results, when the sampling density is $15$ or $10$, the SDF-linearized CBF may lead to a collision behavior, which is illustrated by the red or green lines. When the sampling density is up to $20$, the SDF-linearized CBF, as well as the proposed method, can avoid the collision between vehicles, as seen from the black and blue lines. On the other hand, the computing time of these methods can be obtained by the aid of Simulink/Profiler directly, and they are compared in Tab. II. From the results, one can know that the proposed method is several times faster than the SDF-linearized CBF, no matter the sampling density of the SDF-linearized CBF is chosen as $10,15$ or $20$.
As a brief summary, both methods could avoid the collision, but the proposed method is more efficient than the SDF-linearized CBF, which is a typical example of implicit, optimization-embedded methods in Tab. I.